### v1

@@ Abstract
We develop a deep learning algorithm for approximating functional rational expectations equilibria of dynamic stochastic economies in the sequence space. We use deep neural networks to parameterize equilibrium objects of the economy as a function of truncated histories of exogenous shocks. We train the neural networks to fulfill all equilibrium conditions along simulated paths of the economy. To illustrate the performance of our method, we solve three economies of increasing complexity: the stochastic growth model, a high-dimensional overlapping generations economy with multiple sources of aggregate risk, and finally an economy where households and firms face uninsurable idiosyncratic risk, shocks to aggregate productivity, and shocks to idiosyncratic and aggregate volatility. Furthermore, we show how to design practical neural policy function architectures that guarantee monotonicity of the predicted policies, facilitating the use of the endogenous grid method to simplify parts of our algorithm.
@@
Keywords: deep learning, heterogeneous firms, heterogeneous households, overlapping generations, deep neural networks, global solution method, life-cycle, occasionally binding constraints.
@@ 1 Introduction
In this paper, we bring deep learning solution methods333see, e.g., Azinovic et al., (2022); Maliar et al., (2021); Kahou et al., (2021); Gu et al., (2023); Kase et al., (2023); Gopalakrishna et al., (2024); Valaitis and Villa, (2024); Azinovic and Zemlicka, (2024); Folini et al., (2025); Druedahl and Ropke, (2025) to the sequence space and propose a new global method for computing equilibria in economies with aggregate risk. [^4] We use the expression “sequence space” liberally, referring to the space of truncated sequences of aggregate shocks, not (only) the space of perfect foresight sequences. To be more precise, one could alternatively refer to our approach as moving-average as opposed to sequence-space approach. Exploiting the ergodicity property of a large class of dynamic economies, we approximate the aggregate state vector using a truncated sequence of aggregate shocks. We use deep neural networks to parameterize the mapping from truncated histories of aggregate shocks to equilibrium objects of interest, and optimize them using stochastic gradient descent to minimize equilibrium conditions error over the simulated ergodic set of the economy.
@@
We make four distinct contributions to the literature. First, we show, how deep learning can be used to compute global approximations to equilibria of dynamic stochastic economies in the sequence space. Second, we show, how to construct neural network architectures that predict approximate policy functions, which are guaranteed to satisfy ex-ante known monotonicity and concavity properties. Third, based on our monotonicity-preserving architectures, we show, how to obtain supervised learning targets for policy and price functions, using the endogenous gridpoint method of Carroll, (2006). Fourth, we illustrate the rich applicability and accuracy of our method by solving a challenging overlapping generations model as well as heterogeneous agent economy with heterogeneous firms and households. To the best of our knowledge, we are the first to obtain a global solution to an economy featuring incomplete markets and household heterogeneity in the spirit of Imrohoroğlu, (1989)-Bewley, (1977)-Huggett, (1993)-Aiyagari, (1994)-Krusell and Smith, (1998), together with firm heterogeneity in the spirit of Khan and Thomas, (2008) and Bloom et al., (2018).
@@
As an example, consider the model by Krusell and Smith, (1998) and let $z_{t}$ denote the realization of TFP in period $t$, and let $z^{t}:=(z_{0},z_{1},\dots,z_{t})$. The infinite horizon set-up of the Krusell and Smith, (1998) economy implies the length of the history sequence is also infinite. The state of the economy in a Markov equilibrium in the Krusell and Smith, (1998) economy is the current level of TFP $z_{t}$, together with the cross-sectional distribution of wealth and income. Let $e^{i}_{t}$ denote the idiosyncratic productivity of the household $i$ in period $t$ and let $k_{t}^{i}$ denote their capital holdings. Let $\mu_{t}=\mu_{t}(e,k)$ denote the cross-sectional distribution. Since the distribution, even when discretized into a histogram as in Young, (2010), is a high-dimensional object, this poses a long standing challenge for traditional numeric methods. The approach of Krusell and Smith, (1998) is to approximate the cross-sectional distribution $\mu_{t}$ with a low-dimensional statistic such as aggregate capital $K_{t}$. Instead of approximating the state vector using its moments or a histogram, we approximate the aggregate state vector using sequence-space objects. In particular, we propose to approximate the aggregate state of the economy by a truncated history of exogenous aggregate shocks. In the context of the Krusell and Smith, (1998) model, for example, let $z^{t}_{T}=(z_{t-T},z_{t-T+1},\dots,z_{t-1},z_{t})$ denote the last $T$ realizations of TFP. Instead of computing the equilibrium function $f^{\text{approx}}(z_{t},K_{t})\approx f(z_{t},\mu_{t})$ (or $f^{\text{approx}}(z_{t},\mu_{t}^{\text{histogram}})\approx f(z_{t},\mu_{t})$), we propose computing equilibrium functions $f^{\text{approx}}(z_{t}^{T})\approx f(z_{t},\mu_{t})$. For any $T<\infty$ our approach implies a truncation error. However, in economies featuring ergodic dynamics, this truncation error converges to zero as $T\to\infty$. Hence, we can approximate the information contained in the aggregate state vector arbitrarily well by choosing $T$ sufficiently large. Consequently, our function approximators need to be able to handle large $T$, implying a high-dimensional input $z_{t}^{T}$, potentially even higher dimensional than a histogram of the wealth distribution. We will refer to our approach, which uses a truncated history of aggregate shocks as input as sequence-space approach and to the standard approach (using, for example, the wealth distribution or a lower dimensional sufficient statistic such as standard moments, as input to the equilibrium functions) as state-space approach. We use neural networks to approximate all equilibrium functions that need to be solved for. Like polynomials, neural networks are universal function approximators (Hornik et al.,, 1989). More importantly however, neural networks have shown great promise in ameliorating the curse of dimensionality (Bellman,, 1961) associated high-dimensional function approximation problems, such as solving partial differential equations (see, e.g. Grohs et al.,, 2018; Jentzen et al.,, 2018) or computing equilibria in dynamic economies (see, e.g. Azinovic et al.,, 2022; Maliar et al.,, 2021; Kahou et al.,, 2021; Gu et al.,, 2023). [^5] Section 2 provides a more detailed discussion of the related literature in economics.
@@
The sequence-space formulation has two main advantages in the context of deep learning solution methods. First, the sequence-space formulation is potentially more parsimonious, especially in economies with rich cross-sectional heterogeneity.666e.g. The state-space formulation of the workhorse HANK economy features a state vector that includes a cross-sectional distribution of households over idiosyncratic shocks as well as liquid and illiquid assets. A tensor-product discretization of such a distribution might easily generate tens of thousands of aggregate state variables, whereas a highly accurate sequence-space representation could be obtained using a few hundred of shock lags at quarterly frequency. Second, using truncated shock histories as the only network input mitigates a dangerous feedback loop present in simulation-based deep learning solution methods. While the distribution of shock histories is exogeneous and fixed, [^7] Because the distribution of fundamental shocks is just a primitive of the economic model at hand. the distribution of exogenous states moves during the training as the algorithm updates the approximate policy function. As discussed in Azinovic et al., (2022), a large update policy function might shift the distribution of training states into areas that have been previously only sparsely covered, causing large sample error, and correspondingly large gradients, potentially inducing further jumps in the policy function. While our algorithm still relies on simulating endogenous states an an input for evaluating loss function888e.g. for evaluation of marginal product of capital those do not enter as an input into neural networks parameterizing the core equilibrium objects of the economy.
@@
The second contribution of our paper is that we show how to construct neural network architectures that guarantee monotonicity or even concavity of the resulting approximator with respect to a chosen set of inputs. Rather than directly approximating individual policy function using a neural network that takes the aggregate and idiosyncratic state as input, [^9] As in Azinovic et al., (2022) or Maliar et al., (2021). we build on the operator learning approach of Zhong, (2023). A simple operator network maps the sequence of aggregate shocks to a policy functions over idiosyncratic state variables only. Hence, the neural network predicts different idiosyncratic policy functions for different aggregate states. For the Krusell and Smith, (1998) economy, for example, the neural network predicts a consumption function based on the realized history of aggregate shocks, which takes only individual productivity and asset holdings as input.
This approach allows us to ensure properties of policy functions, such as monotonicity and concavity in the idiosyncratic asset holdings by construction. Being able to guarantee such properties has several advantages. First, in the spirit of economics-inspired neural networks, we encode known properties of the equilibrium policies directly into the architecture of the neural network, so the network does not have learn something which is known ex-ante, rendering the training procedure more robust. Second, in the context of solving for household policies, an always monotone consumption function enables us to apply the method of endogeneous gridpoints of Carroll, (2006), and to train the consumption function in a supervised way, enhancing the stability of the training procedure.
To illustrate the broad applicability of our method, we proceed by applying it to solve three models of increasing complexity. The models we solve include economies featuring multiple discrete as well as continuous continuous shocks aggregate shocks, Markov-switching regimes as well as uncertainty shocks. Furthermore, we show that the approach is compatible with path-following model transformations and market-clearing layers, as introduced in Azinovic and Zemlicka, (2024).
@@ 2 Literature
The idea of approximating the endogenous aggregate state by a truncated sequence of aggregate shocks was first introduced by Chien et al., (2011) in the context of an endowment economy with heterogeneous agents and with a single aggregate shock that takes two discrete values. In their setting, they find that considering the history of the past five aggregate shocks is enough to yield a high-accuracy approximation. In the context of production economies with aggregate risk, Lee, (2025) relies on truncated histories of aggregate shocks to summarize the endogenous aggregate state. The underlying idea of the repeated transition method (RTM) of Lee, (2025) is that in an ergodic economy, all possible states will be realized along a sufficiently long simulation path. Lee, (2025) uses this fact to construct the expected continuation value function for the Bellman equations of the agents in the economy.
@@
Our contribution relative to this literature is generality. While those two papers also develop sequence based global solution methods, the applicability of these methods has limitations that does not pose a significant challenge for our method. Namely, the algorithm of Chien et al., (2011) struggles to resolve economies featuring stronger persistence, such as production economies, because strong persistence implies the need to keep track of long history of shocks. The RTM approach of Lee, (2025) can handle a rich set of environments, including production economies, however, its main bottleneck lies in handling economies with large number of different aggregate shocks. At each iteration the RTM algorithm has to construct the continuation value function for every period. To do so, the algorithm has to split the simulated path into partitions sorted by the realized value of aggregate shock, and then to search through each partition for a realization where the aggregate endogenous state (e.g. wealth distribution) is closest to the next period distribution with respect to some norm. This structure implies that the RTM algorithm works the best in environments with a small number of discrete aggregate shock. Complex stochastic processes that can not be approximated using a parsimonious Markov chain pose a challenge for RTM, as the length of simulated path required for ergodic representation grows with the cardinality of shock discretization. In contrast, we demonstrate that our method can efficiently solve rich economies featuring production and complex driving stochastic processes. Besides that, our algorithm can handle discrete as well as continuous shocks. Furthermore, neural networks can learn from a large number of simultaneous trajectories, allowing for significant acceleration on modern massively parallel computing architectures.
@@
For applications where a local method is sufficient, methods based on linearization in the sequence space have recently become the dominant method of choice for heterogeneous agent models in macroeconomics following the game-changing work by Boppart et al., (2018) and Auclert et al., (2021). Boppart et al., (2018) and Auclert et al., (2021) introduce local methods, which linearize heterogeneous agent models with respect to aggregates, in the sequence space. These methods, together with the accompanying libraries, [^10] See https://github.com/shade-econ/sequence-jacobian for a user friendly library for the method in Auclert et al., (2021). have substantially increased the tractability of heterogeneous agent models with aggregate risk, rendering a wide set of previously intractable models computationally tractable. [^11] See also Reiter, (2009) and Bayer and Luetticke, (2020) for linearization in aggregate variables in the state space. Our method complements this literature by providing a global method, which is hence applicable for models with larger shocks and without certainty equivalence with respect to aggregate risk. The cost of having a global method is that the runtime is substantially longer in comparison to local methods, such as the sequence-space Jacobian method by Auclert et al., (2021).
@@
The toolbox that enables us to approximate functions on a very high dimensional domain of histories of possible economic shocks is deep learning. Deep learning as a tool to compute equilibria in economic models dates back to Duffy and McNelis, (2001), who use a shallow neural network to solve a stochastic growth model with using a parameterized expectations algorithm. Norets, (2012) makes use of neural networks in the context of discrete-state dynamic programming. Closer to this paper Azinovic et al., (2022) and Maliar et al., (2021) use neural networks to approximate equilibrium price and policy functions, and train them to satisfy equilibrium conditions, such as first-order optimality conditions, Bellman equations, and market clearing conditions along the simulated paths of the economy. Azinovic and Zemlicka, (2024) extend these methods by introducing market clearing layers [^12] Market clearing layers are neural network architectures, designed such that an prediction by the neural network is always consistent with market clearing. and step-wise model transformations to robustly solve models with portfolio choice. Kase et al., (2023) show how to use these methods to compute the solutions of economic models for ranges of economic parameters, which are then estimated with maximum likelihood. Han et al., (2022) and Kahou et al., (2021) introduce symmetry preserving neural network architectures and Valaitis and Villa, (2024) use deep learning within a parameterized expectations algorithm. Fernández-Villaverde et al., (2023) use neural networks for a nonlinear forecasting rule within a Krusell and Smith, (1998) algorithm. Kahou et al., (2022) investigate the relationship between transversailty conditions and solutions found by a deep-learning based algorithm. Druedahl and Ropke, (2025) use deep learning to compute optimal choices in a finite horizon lifecycle models with up to eight durable goods and show how deep learning can be applied to models with discrete as well as continuous choices.
@@
Zhong, (2023) introduces operator learning, which we build on in the section 5. While previous deep learning approaches used neural network to directly parameterize mapping from a combination of an aggregate and idiosyncratic state to individual choices, Zhong, (2023) uses neural network to approximate the mapping from an aggregate state into a function which then returns individual choices as a function of individual state. We extend the operator learning method of Zhong, (2023) by showing how to encode monotonicity or concavity of the predicted functions into the network architecture. Sun, 2025a uses neural networks to approximate continuation values in combination with otherwise standard methods. Neural network based solution methods in continuous time are developed or applied in Duarte, (2018), Sauzet, (2021), Gopalakrishna, (2021), Gu et al., (2023), Gopalakrishna et al., (2024) and Payne et al., (2024). [^13] Additional applications of deep learning based solution methods in variety of settings include Folini et al., (2025) and Friedl et al., (2023) (climate economics), Duarte et al., (2021) (partial equilibrium finite horizon model with rich asset choice), Bretscher et al., (2022) (international business cycle), Sun, 2025b (spatial economics), Jungerman, (2023) (monopsony power in the labor market) and Adenbaum et al., (2024) (Bewley, (1977) type economy).
@@
Our contribution relative to the existing literature on deep learning solution methods is twofold. First, we show that deep learning can be efficiently used to construct global approximations to the functional rational expectations equilibria in the sequence space. Second, we construct shape-preserving neural network operator network architectures that allow us to incorporate known shape properties of certain equilibrium functions (e.g. monotonicity and concavity of consumption function in canonical consumption-savings problems) directly into the neural network structure. [^14] The idea of shape-preservation in functional approximation methods for dynamic economies was introduced by Judd and Solnick, (1994), and further developed by Cai and Judd, (2012). The ability to guarantee monotonicity of the consumption function allows us to apply the method of endogenous gridpoints of Carroll, (2006) and hence simplify the solution algorithm for a large class of models featuring standard consumption-savings household problems.
3 Algorithm and first application
To introduce the algorithm we first apply it to the canonical stochastic growth model of Brock and Mirman, (1972). For completeness and clarity the model description is provided section 3.1, and our solution method is introduced in section 3.2.
3.1 Simplest illustrative model

### v2

@@ Abstract
We develop a deep learning algorithm for constructing globally accurate approximations to functional rational expectations equilibria of dynamic stochastic economies in the sequence space. We use deep neural networks to parameterize key equilibrium objects, such as policies or prices, as functions of truncated histories of exogenous shocks. We train the neural networks to satisfy equilibrium conditions along simulated paths of the economy. We illustrate the performance of our method in three environments: (i) a high-dimensional overlapping generations economy with multiple sources of aggregate risk; (ii) an economy with heterogeneous households and firms facing uninsurable idiosyncratic risk and large shocks to idiosyncratic and aggregate volatility; and (iii) a stochastic life-cycle economy with a continuous asset choice and a discrete early-retirement choice that induces local convexities in the continuation values of working-age cohorts. We also propose practical neural policy architectures that guarantee monotonicity of predicted policies, enabling the endogenous grid method to simplify parts of the algorithm. We achieve high precision throughout, with the mean error in equilibrium conditions below $0.2\%$.
@@
Keywords: deep learning, deep neural networks, sequence space, heterogeneous firms, heterogeneous households, overlapping generations, life-cycle, global solution method, discrete choice.
@@ 1 Introduction
We bring deep learning solution methods to the sequence space and propose a new global method for computing equilibria in economies with aggregate risk. [^3] We use the expression “sequence space” liberally, referring to the space of truncated sequences of aggregate shocks, not (only) the space of perfect foresight sequences. To be more precise, one could alternatively refer to our approach as a moving-average rather than a sequence-space approach. Exploiting the ergodicity property of a large class of dynamic economies, we approximate the aggregate state vector using a truncated sequence of aggregate shocks. We use deep neural networks to parameterize the mapping from truncated histories of aggregate shocks to equilibrium objects of interest, and optimize them using stochastic gradient descent to minimize errors in the equilibrium conditions over the simulated ergodic set of the economy.
@@
We make four distinct contributions to the literature. First, we show how deep learning can be used to compute global approximations to equilibria of dynamic stochastic economies in the sequence space. Second, we show how to construct neural network architectures that predict approximate policy functions, which are guaranteed to satisfy ex-ante known monotonicity and concavity properties. Third, based on our monotonicity-preserving architectures, we show how to obtain supervised learning targets for policy and price functions, using the endogenous gridpoint method of Carroll, (2006). Fourth, we illustrate the rich applicability and accuracy of our method by solving three models, each featuring a different computational challenge. We first solve an overlapping generations model with portfolio choice and regime switching. We then proceed to compute a global solution to an economy with a continuum of heterogeneous firms and a continuum of heterogeneous households. To our knowledge, we are the first to do so in an economy featuring incomplete markets and household heterogeneity in the spirit of Imrohoroğlu, (1989)-Bewley, (1977)-Huggett, (1993)-Aiyagari, (1994)-Krusell and Smith, (1998), together with firm heterogeneity in the spirit of Bloom et al., (2018) and Khan and Thomas, (2008). Lastly, we apply our method to solve an overlapping generations model with intra- and intergenerational risk and a discrete early-retirement choice. The discrete choices lead to local convexities in the continuation values of working-age cohorts, rendering their Euler equations insufficient for uniquely characterizing their optimal savings policies.
@@
To describe the main idea of our method, we now take the canonical Krusell and Smith, (1998) model as an example. Let $z_{t}$ denote the realization of stochastic TFP in period $t$, and let $z^{t}:=(z_{0},z_{1},\dots,z_{t})$ denote the history of realized shocks. In an infinite horizon model the history of shocks becomes infinitely long. The state of the economy is the current level of TFP $z_{t}$, together with the cross-sectional distribution of households over wealth and productivity. Let $e^{i}_{t}$ denote the idiosyncratic productivity of the household $i$ in period $t$, let $k_{t}^{i}$ denote their capital holdings, and let $\mu_{t}=\mu_{t}(e,k)$ denote the cross-sectional distribution. Since the distribution, even when discretized into a histogram as in Young, (2010), is a high-dimensional object, the presence of the distribution in the state vector poses a long-standing challenge for solution methods. [^4] See, for example, Moll, (2025). The approach of Krusell and Smith, (1998) is to approximate the cross-sectional distribution $\mu_{t}$ with a low-dimensional statistic such as aggregate capital $K_{t}$. Instead of approximating the state vector using its moments or a histogram, we build on the seminal perturbation approaches in the sequence space developed in Boppart et al., (2018) and Auclert et al., (2021) and approximate the aggregate state vector using sequence-space objects. In particular, we propose to approximate the aggregate state of the economy by a truncated history of exogenous aggregate shocks. In the context of the Krusell and Smith, (1998) model, for example, let $z^{t}_{T}=(z_{t-T},z_{t-T+1},\dots,z_{t-1},z_{t})$ denote the last $T$ realizations of TFP. Instead of solving for equilibrium functions $f^{\text{approx}}(z_{t},K_{t})\approx f(z_{t},\mu_{t})$, we propose to solve for equilibrium functions $f^{\text{approx}}(z^{t}_{T})\approx f(z_{t},\mu_{t})$. For any $T<\infty$ our approach implies a truncation error. However, in economies featuring ergodic dynamics, this truncation error converges to zero as $T\to\infty$. Hence, we can approximate the information contained in the aggregate state vector arbitrarily well by choosing $T$ sufficiently large. Consequently, our function approximators need to be able to handle large $T$, implying a high-dimensional input $z_{t}^{T}$, potentially even higher dimensional than a histogram of the wealth distribution. We will refer to our approach, which uses a truncated history of aggregate shocks as input as sequence-space approach and to the standard approach (using, for example, the wealth distribution or a lower dimensional sufficient statistic such as standard moments, as input to the equilibrium functions) as state-space approach.
@@
We use neural networks to approximate all equilibrium functions that need to be solved for. Like polynomials, neural networks are universal function approximators (Hornik et al.,, 1989). More importantly however, neural networks have shown great promise in ameliorating the curse of dimensionality (Bellman,, 1961) associated with high-dimensional function-approximation problems, such as solving partial differential equations (see, e.g. Grohs et al.,, 2018; Jentzen et al.,, 2018) or computing equilibria in dynamic economies (see, e.g. Azinovic et al.,, 2022; Maliar et al.,, 2021; Kahou et al.,, 2021; Gu et al.,, 2023). [^5] Section 2 provides a more detailed discussion of the related literature in economics. The sequence-space formulation has two main advantages in the context of deep learning solution methods. First, the sequence-space formulation is potentially more parsimonious, especially in economies with rich cross-sectional heterogeneity.666e.g. The state-space formulation of the workhorse HANK economy features a state vector that includes a cross-sectional distribution of households over idiosyncratic shocks as well as liquid and illiquid assets. A tensor-product discretization of such a distribution might easily generate tens of thousands of aggregate state variables, whereas a highly accurate sequence-space representation could be obtained using a few hundred of shock lags at quarterly frequency. Second, using truncated shock histories as the only network input mitigates a dangerous feedback loop present in simulation-based deep learning solution methods. While the distribution of shock histories is exogenous and fixed, [^7] Because the distribution of fundamental shocks is just a primitive of the economic model at hand. the distribution of endogenous states moves during the training as the algorithm updates the approximate policy function. As discussed in Azinovic et al., (2022), a large update to the policy function might shift the distribution of training states into areas that have been previously only sparsely covered, causing large sample error, and correspondingly large gradients, potentially inducing further jumps in the policy function. While our algorithm still relies on simulating endogenous states as an input for evaluating loss function888e.g. for evaluation of marginal product of capital those do not enter as an input into neural networks parameterizing the core equilibrium objects of the economy.
@@
The second contribution of our paper is that we show how to construct neural network architectures that guarantee monotonicity or even concavity of the resulting approximator with respect to a chosen set of inputs. Rather than directly approximating individual policy function using a neural network that takes the aggregate and idiosyncratic state as input, [^9] As in Azinovic et al., (2022) or Maliar et al., (2021). we build on the operator learning approach of Zhong, (2023). An operator network maps the sequence of aggregate shocks to policy functions. The predicted policy functions then map idiosyncratic state variables into the individual policy choices. Therefore, instead of having to repeatedly evaluate the neural network for many combinations of aggregate and idiosyncratic states, the neural network predicts a single policy function for each aggregate state. The predicted policy function is then evaluated for all idiosyncratic states of interest. For the Krusell and Smith, (1998) economy, for example, the neural network predicts a consumption function based on the realized history of aggregate shocks, which then takes only individual productivity and asset holdings as input. Additionally, ensuring monotonicity allows us to implement parts of the algorithm using the endogenous grid method (EGM) (Carroll,, 2006). The use of EGM is particularly advantageous when tackling models where agents face a mix of discrete and continuous choices since it enables the use of an upper envelope algorithm (Druedahl and Jørgensen,, 2017; Dobrescu and Shanker,, 2023).
@@ 2 Literature
The idea of approximating the endogenous aggregate state by a truncated sequence of aggregate shocks was first introduced by Chien et al., (2011) in the context of an endowment economy with heterogeneous agents and with a single aggregate shock that takes two discrete values. In their setting, they find that considering the history of the past five aggregate shocks is enough to yield a high-accuracy approximation. In the context of production economies with aggregate risk Lee, (2025) introduces the repeated transition method (RTM). The underlying idea of the RTM is that, in an ergodic economy, all possible states will be realized along a sufficiently long simulation path. Lee, (2025) uses this fact to construct the continuation value function for the Bellman equations of the agents in the economy. The RTM iterates over simulating a long path of the economy. In each iteration it constructs the continuation values by using the previously simulated path as a look-up table. Our contribution relative to this literature is generality. While those two papers also develop sequence-based global solution methods, the applicability of these methods has limitations that do not pose a significant challenge for our method. Namely, the algorithm of Chien et al., (2011) struggles to resolve economies featuring stronger persistence, such as production economies, because strong persistence implies the need to keep track of a long history of shocks. The RTM approach of Lee, (2025) can handle a rich set of environments, including production economies; however, its main bottleneck lies in handling economies with a large number of different aggregate shocks. At each iteration the RTM algorithm has to construct the continuation value function for every period. To do so, the algorithm has to split the simulated path into partitions sorted by the realized value of aggregate shock, and then to search through each partition for a realization where the aggregate endogenous state (e.g. wealth distribution) is closest to the next period distribution with respect to some norm. This structure implies that the RTM algorithm works best in environments with a small number of discrete aggregate shocks. Complex stochastic processes that cannot be approximated using a parsimonious Markov chain pose a challenge for RTM, as the length of simulated path required for ergodic representation grows with the cardinality of shock discretization. In contrast, we demonstrate that our method can efficiently solve rich economies featuring production and complex driving stochastic processes. Besides that, our algorithm can handle discrete as well as continuous shocks. Furthermore, neural networks can learn from a large number of simultaneous trajectories, allowing for significant acceleration on modern massively parallel computing architectures. [^10] Global solution methods to solve models on the simulated paths using state-space approaches include Brumm and Hußmann, (2025), Levintal, (2018), and Azinovic et al., (2023)
Instead of using the truncated history of exogenous shocks as a summary statistic for the aggregate state, recent work by Yang et al., (2025) and Guarda, (2025) uses short histories of endogenous aggregate variables in the spirit of Krusell and Smith, (1998). Yokomoto, (2025) implements the Krusell and Smith, (1998) algorithm using neural networks as function approximators. Yang et al., (2025) and Yokomoto, (2025) use prices as their summary statistic for the endogenous aggregate state, allowing for fast computation of market clearing prices in the simulation step. Yang et al., (2025) update policies with a policy gradient algorithm to maximize simulated lifetime rewards, while Yokomoto, (2025) updates policies using the Bellman equation. Households’ decisions depend on the endogenous aggregate state only through prices. Hence, households directly track the aggregate variable that enters into their optimization problem. In Yang et al., (2025) very short histories are sufficient to forecast conditional future prices well, keeping the state-space for households low-dimensional and thereby addressing the challenges raised in Moll, (2025). [^11] Applications in Yang et al., (2025) include the Krusell and Smith, (1998) model, a Huggett, (1993) model with aggregate risk, and a one asset HANK model along the lines of Auclert et al., (2021). Although we focus on histories of exogenous shocks, our method can easily be extended to additionally use histories of endogenous aggregate variables, such as prices. In contrast to Yang et al., (2025) our approach works directly with equilibrium conditions and hence allows us to obtain supervised learning targets for household policies. Furthermore, our approach applies to models where the individual problem depends directly on a high dimensional aggregate state and not only through a few low-dimensional aggregate statistics, such as prices. [^12] See, for example, Payne et al., (2024).
@@
Perturbation methods in the sequence space have become the method of choice for solving heterogeneous agent models following the game-changing work by Boppart et al., (2018) and Auclert et al., (2021). Boppart et al., (2018) and Auclert et al., (2021) introduce local methods, which linearize heterogeneous agent models with respect to aggregates, in the sequence space. [^13] See also Reiter, (2009) and Bayer and Luetticke, (2020) for linearization in aggregate variables in the state space. These methods, together with the accompanying libraries, [^14] See https://github.com/shade-econ/sequence-jacobian for a user friendly library for the method in Auclert et al., (2021). have substantially increased the tractability of heterogeneous agent models with aggregate risk, rendering a wide set of previously intractable models computationally tractable. Our method complements this literature by providing a global method, which is hence applicable for models with larger shocks and without certainty equivalence with respect to aggregate risk. The cost of having a global method is that the runtime is substantially longer in comparison to local methods, such as the sequence-space Jacobian method by Auclert et al., (2021).
@@
The toolbox that enables us to approximate functions on the very high dimensional domain of long histories of economic shocks is deep learning. Deep learning as a tool to compute equilibria in economic models dates back to Duffy and McNelis, (2001), who use a neural network to solve a stochastic growth model using a parameterized expectations algorithm. Norets, (2012) makes use of neural networks in the context of discrete-state dynamic programming. Closer to this paper Azinovic et al., (2022) and Maliar et al., (2021) use neural networks to approximate equilibrium price and policy functions, and train them to satisfy equilibrium conditions, such as first-order optimality conditions, Bellman equations, and market clearing conditions along the simulated paths of the economy. Azinovic and Zemlicka, (2024) extend these methods by introducing market clearing layers [^15] Market clearing layers are neural network architectures, designed such that a prediction by the neural network is always consistent with market clearing. and step-wise model transformations to robustly solve models with portfolio choice. Kase et al., (2023) show how to use deep learning to compute the solutions of economic models for ranges of economic parameters, speeding up the maximum likelihood estimation of these parameters. Han et al., (2022) and Kahou et al., (2021) introduce symmetry preserving neural network architectures and Valaitis and Villa, (2024) use deep learning within a parameterized expectations algorithm. Fernández-Villaverde et al., (2023) use neural networks for a nonlinear forecasting rule within a Krusell and Smith, (1998) algorithm. Kahou et al., (2022) investigate the relationship between transversality conditions and solutions found by a deep-learning based algorithm. Druedahl and Ropke, (2026) use deep learning to compute optimal choices in finite-horizon life-cycle models with up to eight durable goods. Druedahl et al., 2026a show how the methodology of Druedahl and Ropke, (2026) can be applied to solve non-convex life-cycle models featuring a mixture of discrete and continuous choices. We view our discrete-choice algorithm introduced in section 6 as complementary to Druedahl et al., 2026a : we focus on general equilibrium setups while keeping the optimization problem of an individual agent comparatively simple, whereas Druedahl et al., 2026a solve complex individual optimization problems with many states and choices in partial equilibrium. In recent work Druedahl et al., 2026b build on our approach to construct a global solution to a canonical HANK economy with aggregate risk using deep learning in the sequence space. They study the effect of aggregate uncertainty on fiscal multipliers and impulse responses to aggregate shocks.
@@
Zhong, (2023) introduces operator learning, which we build on in section 5. While previous deep learning approaches used neural networks to directly parameterize the mapping from a combination of an aggregate and idiosyncratic state to individual choices, Zhong, (2023) uses neural networks to approximate the mapping from an aggregate state into a function which then returns individual choices as a function of individual state. We extend the operator learning method of Zhong, (2023) by showing how to encode monotonicity or concavity of the predicted functions into the network architecture. Sun, 2025a uses neural networks to approximate continuation values in combination with otherwise standard methods. Neural network based solution methods in continuous time are developed or applied in Duarte, (2018), Sauzet, (2021), Gopalakrishna, (2021), Gu et al., (2023), Gopalakrishna et al., (2024) and Payne et al., (2024). [^16] Additional applications of deep learning based solution methods in variety of settings include Folini et al., (2025) and Friedl et al., (2023) (climate economics), Carvalho et al., (2025) (production networks), Kase et al., (2025) (HANK with financial frictions), Duarte et al., (2021) (partial equilibrium finite horizon model with rich asset choice), Bretscher et al., (2022) (international business cycle), Sun, 2025b (spatial economics), Jungerman, (2023) (monopsony power in the labor market), Adenbaum et al., (2024) (Bewley, (1977) type economy), and Reiter, (2025) (hybrid solution approach for OLG).
@@
Our contribution relative to the existing literature on deep learning solution methods is twofold. First, we show that deep learning can be efficiently used to construct global approximations to the functional rational expectations equilibria in the sequence space. Second, we construct shape-preserving neural network operator network architectures that allow us to incorporate known shape properties of certain equilibrium functions (e.g. monotonicity and concavity of consumption function in canonical consumption-savings problems) directly into the neural network structure. [^17] The idea of shape-preservation in functional approximation methods for dynamic economies was introduced by Judd and Solnick, (1994), and further developed by Cai and Judd, (2012). The ability to guarantee monotonicity of the consumption function allows us to apply the method of endogenous gridpoints of Carroll, (2006) and hence simplify the solution algorithm for a large class of problems featuring optimization problems that are amenable to EGM.
3 Algorithm and illustrative application
3.1 Model